1) A machinery upgrade project will cost $14 million, and produce cost savings of $1 million in perpetuity.

What is the project NPV if the required rate of return is 11%?

Enter answer in millions, to two decimal places.

2)

Your company is considering a project that requires an initial investment of $12 million, and is expected to produce cash flows of $4 million each year for 10 years. At the end of year 11, the project will require site cleanup that is expected to cost $11 million. What is the project NPV if the company uses 12% as its cost of capital?

Enter answer in millions, to two decimal places.

3)A project with an initial cost of $591 is expected to produce a stable cash inflow for 9 years. Using a required return of 11%, it has been determined that the project's NPV is 82. What must be the size of each of the annual cash flows?
Answer #1

Please provide rating..

Answer 1
Machine cost 14 mil
Annual cost saving 1 mil
required rate = 11%
PV of annuity = A/I = 1mil/11%                  9.09 mil
NPV = PV of cash flow - initial cost =
=9.09-11                 (4.91) mil
Note : NPV is negative
Answer 2
i ii iii=i*ii
Year Cash flow PVIF @ 12% present value
0 -12              1.0000              (12.00)
1 4              0.8929                  3.57
2 4              0.7972                  3.19
3 4              0.7118                  2.85
4 4              0.6355                  2.54
5 4              0.5674                  2.27
6 4              0.5066                  2.03
7 4              0.4523                  1.81
8 4              0.4039                  1.62
9 4              0.3606                  1.44
10 4              0.3220                  1.29
11 -11              0.2875                (3.16)
                 7.44
NPV =              7.44 mil
Answer 3
Lets assume that annual cash flow = x
Therefore we have PV = 82 at 11% required rate of return
PVIFA @ 11% for 9 year =          5.5370 (From PVIFA table)
we have below equation for NPV
X*5.5370-591 = 82
X = =(591+82)/5.537
Annual cash flow          121.55